john is three times the age of melissa. six years ago, the sum of their ages was 48. how old are they now?
Answer by ikleyn(52792)

.
"six years ago, the sum of their ages was 48" - hence, now the sum of their ages is 48 + 2*6 = 48 + 12 = 60 years.


So, we know that J is three times M and the sum J + M is 60.


Hence (solving mentally), M  is 15 years old, while John is 3*15 = 45 years old now.    ANSWER